OpenStudy (ccastudentgrade10):

fan and medal pls help!! simplify sqrt9x^3 * sqrt18x and (27)^-2/3 i have multiple problems like this on the worksheet, so pls explain your steps if you can! i missed a week of alg2 and im lost

OpenStudy (anonymous):

\[\sqrt{9x}.\sqrt{18x} = \sqrt{9.18.x.x} = \sqrt{3.3.3.3.2.x.x} = ???\]

OpenStudy (anonymous):

\[26^{\frac{-2}{3}} = \frac{1}{27^{\frac{2}{3}}} = \frac{1}{\sqrt[3]{27*27}} = \frac{1}{\sqrt[3]{3.3.3.3.3.3}} = ??\]

OpenStudy (ccastudentgrade10):

i don't understand what the answer is??

OpenStudy (anonymous):

\[\sqrt[3]{a.a.a} = a\]

OpenStudy (ccastudentgrade10):

and didn't you forget the ^3 in the first problem?

OpenStudy (anonymous):

Yep yes I forgot.

OpenStudy (ccastudentgrade10):

and why is it 26, not 27 in the second problem?

OpenStudy (anonymous):

insted of x.x it will become x.x.x.x

OpenStudy (ccastudentgrade10):

so is this answer to the second one 3?

OpenStudy (anonymous):

first 26 is typo it should be 27

OpenStudy (ccastudentgrade10):

ohhh okay so the answer is 3?

OpenStudy (anonymous):

There are total of 6's 3 so it should be 3.3 = 9 so 1 / 9.

OpenStudy (ccastudentgrade10):

okay thank you! could you possibly redo the first problem with the ^3?

OpenStudy (anonymous):

\[\sqrt{9.9.x.x.x.x} = ??\]

OpenStudy (ccastudentgrade10):

i don't understand at all can you start from the beginning? @baljeet16

OpenStudy (ccastudentgrade10):

@AravindG ??

OpenStudy (anonymous):

Yes sure

OpenStudy (anonymous):

\[\sqrt{9x^{3}}\sqrt{18x} = \sqrt{9.x^3.18.x} = \sqrt{9.9.2.x^2.x^2} = ??\]

OpenStudy (ccastudentgrade10):

so is the answer 4x/18 ?? @baljeet16

OpenStudy (anonymous):

where did you got 4 ??

OpenStudy (ccastudentgrade10):

the 4 x's

OpenStudy (anonymous):

\[\sqrt{a.a} = a \\ \sqrt{a^2.a^2} = a^2\]

OpenStudy (ccastudentgrade10):

x/9 ??

OpenStudy (anonymous):

\[9x.\sqrt{2}\]

OpenStudy (ccastudentgrade10):

OH okay

OpenStudy (ccastudentgrade10):

could you help me w two more problems??

OpenStudy (anonymous):

yes sure.

OpenStudy (ccastudentgrade10):

simplify and rationalize denominators: 1. 6/8-sqrt3 2. sqrtw^4/a^3

OpenStudy (ccastudentgrade10):

@baljeet16

OpenStudy (anonymous):

yes.

OpenStudy (ccastudentgrade10):

what? the questions are above @baljeet16

OpenStudy (anonymous):

question 1 is not clear.

OpenStudy (ccastudentgrade10):

okay, i need to simplify the fraction 6/8 - sqrt 3

OpenStudy (ccastudentgrade10):

and rationalize the denominator

OpenStudy (ccastudentgrade10):

6/(8- sqrt3) is that better?

OpenStudy (anonymous):

yes

OpenStudy (ccastudentgrade10):

okay, so what do i do?

OpenStudy (anonymous):

\[\frac{6}{8 - \sqrt{3}} = \frac{6}{8 - 1.732} = ????\] Can you use calculator.

OpenStudy (ccastudentgrade10):

no, no calculators :/

OpenStudy (ccastudentgrade10):

8-1.732 is 6.268 @baljeet16

OpenStudy (ccastudentgrade10):

so 6/6.268?

OpenStudy (anonymous):

\[\sqrt{\frac{a^4}{x^3}} = \sqrt{\frac{a^2.a^2}{x^2.x}} = \frac{a}{x}.x^{\frac{-1}{2}}\]

OpenStudy (anonymous):

yes

OpenStudy (ccastudentgrade10):

is that the final answer? (for the second one)

OpenStudy (anonymous):

Can you understand the second explaination.

OpenStudy (ccastudentgrade10):

yes i understand what you did, but is that the final answer? @baljeet16

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

No its little wrong.

OpenStudy (anonymous):

can you find whats wrong?

OpenStudy (anonymous):

in second question?

OpenStudy (ccastudentgrade10):

should the answe be a/x^2 instead of a/x ? @baljeet16

OpenStudy (ccastudentgrade10):

or should the 1/2 be positive? @baljeet16

OpenStudy (anonymous):

it should be\[\frac{a^2}{x}.x^{\frac{-1}{2}}\]

OpenStudy (ccastudentgrade10):

ohhh i see what you did @baljeet16

OpenStudy (ccastudentgrade10):

could you possibly help me with one last problem @baljeet16 ?? if not its okay uve helped me a ton

OpenStudy (anonymous):

Okay ask.

OpenStudy (ccastudentgrade10):

simplify 6sqrt40 - 2sqrt90 - 3sqrt160

OpenStudy (ccastudentgrade10):

@baljeet16

OpenStudy (anonymous):

\[6.\sqrt{2.2.10} - 2.\sqrt{3.3.10} - 3.\sqrt{4.4.10} = 6.2.\sqrt{10} - 2.3.\sqrt{10} - 3.4.\sqrt{10} = ???\]

OpenStudy (ccastudentgrade10):

5sqrt-10 ?

OpenStudy (ccastudentgrade10):

@baljeet16

OpenStudy (anonymous):

-5 sqrt(10) yes

OpenStudy (ccastudentgrade10):

thank you!

OpenStudy (anonymous):

sorry its \[-6.\sqrt{10}\]

OpenStudy (ccastudentgrade10):

okay, thanks!